OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, −4) and a directrix of y = 4.

OpenStudy (anonymous):

I miss this stuff :D Got any idea what focus and directrix mean?

OpenStudy (anonymous):

To be honest, not a clue.

OpenStudy (solomonzelman):

If you can tell me the equation of this parabola, so that we can start on the calculus part of your problem.

OpenStudy (anonymous):

That's the whole question.

OpenStudy (solomonzelman):

I can see what you are asked, but I tried to nicely ask you to find the equation of the parabola first.

OpenStudy (anonymous):

I'm sorry. I don't know how to do that either. This is my weak point in math.

OpenStudy (solomonzelman):

Okay, you know it is a vertical parabola, since it's directrix is "y="

OpenStudy (anonymous):

...great. The OP left. @SolomonZelman You basically asked the OP to do the question himself :/

OpenStudy (solomonzelman):

@Supreme_Kurt maybe, if you don't know how to help the user, you can try not to interrupt? That would be great.

OpenStudy (solomonzelman):

Okay, mdf, So, your focus is (0,-4), and by "0" the x coordinate of the minimum point is given. Now, the directrix, gives it's y-value, and in this case it is y=4. So you know that the minimum point is (0,4) \(\large\color{black}{ (x-h)^2=4p(y-k) }\) would be essentially giving the equation of the parabola. You know that the focus is 8 units away from the vertex.

OpenStudy (solomonzelman):

\(\large\color{black}{ (x-0)^2=4(8)(y-4) }\) See how I am plugging it in?

OpenStudy (solomonzelman):

\(\large\color{black}{ x^2=32y-128 }\) \(\large\color{black}{ 32y=- x^2-128 }\) \(\large\color{black}{ y=-\frac{1}{32}x^2-4 }\)

OpenStudy (solomonzelman):

To derive it, you shall use just the power rule.

OpenStudy (anonymous):

Good thing the OP has left. Your answer is wrong. Care to try again?

OpenStudy (solomonzelman):

yes, I would want to remember how to do it correctly-:(

OpenStudy (solomonzelman):

I'll try on my paper.

OpenStudy (anonymous):

First, why don't you apologize to your former tutor? :P

OpenStudy (anonymous):

Anyway, you're right in essence. the equation of a vertical parabola is \[\Large (x-h)^2 = 4p(y-k)\] with (h,k) being the vertex.

OpenStudy (anonymous):

First problem: Wrong vertex. care to fix that? What's the correct vertex?

OpenStudy (solomonzelman):

thew vertex is between.

OpenStudy (solomonzelman):

(0,0) so I would just plug in 4, and it would be \(\large\color{black}{(x-0)^2=4(4)(y-0) }\) \(\large\color{black}{x^2=16y }\) \(\large\color{black}{\frac{1}{16}x^2=y }\)

OpenStudy (anonymous):

Still no.

OpenStudy (anonymous):

Why would you plug in p = 4?

OpenStudy (solomonzelman):

the distance between vertex and focus is 4?

OpenStudy (solomonzelman):

yeah... math was never my thing. I'll watch a video and be back I guess.

OpenStudy (anonymous):

Well, you could put it that way. But I prefer to think of p as the difference between the vertex y-value and the focus y-value (for vertical parabolas). In that case, p=-4 That way, you can immediately see that it's a downward-facing parabola with a MAXIMUM rather than a minimum, as you stated...

OpenStudy (solomonzelman):

:) it's over -16.

OpenStudy (solomonzelman):

I truely suck

OpenStudy (solomonzelman):

truly

OpenStudy (anonymous):

No, you don't :/ You just need practice. And humility. Never forget humility, ok? ^^

OpenStudy (solomonzelman):

truly

OpenStudy (solomonzelman):

EVeryone sucks in his own way, or at least anyone. But I think I learned again. I watched that video with -p